# CLOVA: A Closed-LOop Visual Assistant with Tool Usage and Update

### v1

@@ Abstract
Leveraging large language models (LLMs) to integrate off-the-shelf tools (e.g., visual models and image processing functions) is a promising research direction to build powerful visual assistants for solving diverse visual tasks. However, the learning capability is rarely explored in existing methods, as they freeze the used tools after deployment, thereby limiting the generalization to new environments requiring specific knowledge. In this paper, we propose CLOVA, a Closed-LOop Visual Assistant to address this limitation, which encompasses inference, reflection, and learning phases in a closed-loop framework. During inference, LLMs generate programs and execute corresponding tools to accomplish given tasks. The reflection phase introduces a multimodal global-local reflection scheme to analyze whether and which tool needs to be updated based on environmental feedback. Lastly, the learning phase uses three flexible manners to collect training data in real-time and introduces a novel prompt tuning scheme to update the tools, enabling CLOVA to efficiently learn new knowledge without human involvement. Experimental results show that CLOVA outperforms tool-usage methods by 5% in visual question answering and multiple-image reasoning tasks, by 10% in knowledge tagging tasks, and by 20% in image editing tasks, highlighting significance of the learning capability for general visual assistants.
Figure 1: CLOVA is a general visual assistant that updates both LLMs and visual models via inference, reflection, and learning in a closed-loop framework. During inference, CLOVA uses LLMs to integrate visual tools to accomplish given tasks. In reflection, CLOVA identifies models that require updating based on environmental feedback. Finally, in learning, CLOVA collects data and updates models accordingly.
footnotetext: ✉  Corresponding author: Qing Li.
@@ 1 Introduction
The progresses of large language models (LLMs) [60, 38, 7, 74] and various visual models [47, 48, 6, 36, 26] offer feasible ways to build general visual assistants. Given a task with language instruction, a prevalent way is to leverage LLMs to generate programs, and then execute off-the-shelf visual tools (e.g., OpenCV functions and visual models) to accomplish the task according to the generated program [11, 8, 32, 58, 53, 66, 65]. For example, when asking the question “What is the person to the left of the umbrella doing?”, a potential solution entails LLMs sequentially (1) using the detection model to locate the umbrella, (2) cropping the image region left to the umbrella, (3) using the detection model to locate the person, and (4) finally querying a visual question answering (VQA) model with “What is the person doing?”. Due to the remarkable compositionality of this process, tool-usage methods achieve impressive performance in solving complicated visual tasks, such as compositional VQA, vision-and-language navigation, and language-guided image editing [72, 73, 30, 67].
However, the learning capability is rarely explored in existing tool-usage methods. Most of them simply freeze used tools after deployment, limiting the generalization to new environments that require specific knowledge, e.g., how to tackle problems with visual concepts and environments that are novel to the tools. We’ve demonstrated a few in Fig. 1. For example, the user might instruct the visual assistant to tag the face of the movie director Bong Joon-ho in a picture. However, the face classification tool invoked by the assistant might not be familiar to him, and ultimately the assistant returns with a faulty response. In such cases, we may expect the assistant to be able to learn this missing piece of knowledge on celebrity after the initial failure and generalize it to other tasks, e.g. tagging Bong Joon-ho in another picture. Thus, it is crucial to endow visual assistants with the learning capability, enabling them to quickly adapt to new tasks and environments.
@@
In this paper, we propose CLOVA, a Closed-LOop Visual Assistant that updates used tools via closed-loop learning to better adapt to new environments, as shown in Fig. 1. CLOVA comprises three phases: inference, reflection, and learning. During inference, CLOVA employs LLMs to generate programs and execute corresponding tools to accomplish the task at hand. Subsequently in the reflection phase, CLOVA uses environmental feedback to produce critiques, identifying tools that need to be updated. Finally, the learning phase autonomously collects data and updates tools accordingly. In this case, CLOVA enables tools to be updated autonomously, continually improving their ability to adapt to diverse environments.
@@
To establish such a closed-loop learning framework, we must overcome three key challenges. 1) How to identify tools that need to be updated. The multi-step nature of generated programs and the diverse errors in them make identifying the faulty tools challenging. 2) How to collect training data. Real-time collection is necessary, since the knowledge that needs to be learned is unpredictable, causing the difficulty. 3) How to efficiently update tools amid their scale and the quality of collected data. The tools used by the assistant are generally backed by large models and inefficient to be updated, while naive fine-tuning could also lead to unacceptable catastrophic forgetting [34]. The noise within the collected data further perplexes the training issue.
@@
We propose several techniques to tackle these challenges. First, we introduce a multimodal global-local reflection scheme, which resorts to LLMs to identify tools that need to be updated from both global and local aspects. For the second challenge, three data collection manners are employed, including inferring answers by LLMs, searching on the Internet, and searching from open-vocabulary datasets. Lastly, we develop a training-validation prompt tuning scheme for the tools, which includes instance-wise prompt tuning and a subsequent prompt validation stage, where learned prompts that fail to predict the validation data will be discarded. The learning phase also updates LLMs by storing correct examples and incorrect examples with critiques as in-context examples, which will be used in future inference. As a result, CLOVA efficiently updates tools in challenging environments with noisy data, while avoiding existing knowledge being wiped out.
@@
We apply CLOVA to compositional VQA and multiple-image reasoning tasks, using the GQA [17] and NLVRv2 [57] datasets. Additionally, we manually collect data for image editing and factual knowledge tagging tasks. CLOVA outperforms existing tool-usage methods by 5% in compositional visual question answering and multiple-image reasoning tasks, by 10% in knowledge tagging tasks, and by 20% in image editing tasks, showing the significance of the learning capability for general visual assistants.
In summary, our contributions are four-fold. 1) We build CLOVA, a visual assistant that updates its tools within a closed-loop learning framework for better adaptation to new environments. 2) We propose a multimodal global-local reflection scheme, capable of identifying tools in need of updates. 3) We employ three flexible manners to collect training data in real-time, without any human involvement. 4) We introduce a training-validation prompt tuning scheme, enabling models to update themselves in challenging environments efficiently while preserving existing knowledge.
@@ 2.1 General Visual Assistant
Benefiting from the advancements of LLMs [60, 38, 7, 74] and visual models [47, 23, 36, 48], visual assistants have achieved great progresses. Some methods concatenate and train LLMs with visual models in an end-to-end manner, where representative work includes LLaVA [28], Otter [24], MMICL [76], Kosmos-2 [43], and Flamingo [1], etc. In addition, some work extends the idea of tool usage for AI assistants from natural language processing [46, 40, 4, 45, 51, 12] to computer vision. By providing in-context examples, VISPROG [11] and ViperGPT [58] generate programs to use visual tools. Following this idea, some work improves performance by collecting instruction-following data [72, 29, 42], adding more tools [53, 27], and designing more dedicated tool-usage procedures [68, 73, 31, 32, 67, 30, 13]. The most related work to CLOVA is AssistGPT [8] and OpenAGI [9]. The two methods update LLMs after development through in-context learning and reinforcement learning, respectively. Different from them, CLOVA has the ability to update both LLMs and visual models through its reflection and learning phases. This allows CLOVA to better adapt to new environments. In addition, the closed-loop framework enables us to set a separate training stage, going beyond zero-shot or few-shot visual assistants. Comparisons between CLOVA and some representative tool-usage methods are shown in Tab. 1.
Method Visual Tool Reflection Update LLMs Update VMs
@@
Table 1: Comparisons with representative tool-usage methods, where VMs means visual models.
@@ 2.2 Reflection of LLMs
Reflection has become a remedy in case LLMs cannot generate good responses in a single attempt [39, 71, 41, 5, 22]. Reflection methods send outputs back to LLMs to obtain critiques and further improve the outputs. These critiques take the form of scores or natural language [35, 55, 37, 79]. To generate better critiques, some methods employ instruction tuning [70, 49] or reinforcement learning [45, 3]. Recently, Huang et al. [15] revealed that LLMs struggle to provide accurate critiques for complex tasks. One way to address this issue is incorporating external information such as human-desired results into LLMs [71, 69, 2]. Unlike existing methods that rely solely on feedback in the language modality, our method generates reflection using all multimodal intermediate results. In addition, our method incorporates both global and local aspects for reflection, instead of only the global aspect. These bring more effective critiques for compositional tasks.
2.3 Prompt Tuning
Prompt tuning is an efficient technique to update neural networks, achieving impressive performance in both NLP [54, 19] and computer vision [80, 52, 81, 16]. Hard prompt tuning and soft prompt tuning are two kinds of commonly used methods. Hard prompt tuning develops interpretable tokens (e.g., texts and image regions) to guide model prediction, which are usually obtained by manually designing [59, 47], retrieval [75], and model generation [14, 44]. Soft prompt tuning learns vectors as prompts via gradient-based optimization. VPT [18] and CoOp [78] learn prompts for vision encoders and text encoders, respectively. To handle diverse data, CoCoOp [77] learns to generate prompts for unknown classes, ProDA [33] builds a Gaussian distribution for prompts, and MaPLe [21] uses both the text and visual prompts. In addition, some methods employ soft prompts for continual learning, which learn prompts for different classes and produce adaptive prompts during inference. Representative methods include PIVOT [62], DualPrompt [63], and L2P [64]. Different from existing methods, our training-validation prompt tuning scheme discards harmful prompts, leading to more stable learning processes when the quality of training data is subpar.
@@ 3.1 Overview
CLOVA has three phases: inference, reflection, and learning, as shown in Fig. 2. In the inference phase, CLOVA uses LLMs to generate programs and executes corresponding tools to solve the task. The reflection phase introduces a multimodal global-local reflection scheme that uses LLMs to generate critiques, identifying which tool needs to be updated. During learning, we employ three manners to collect training data and use a training-validation prompt tuning scheme to update the tools.
@@
Table 2: Used tools in CLOVA, categorized based on whether the tool is updated in our method or not. Details on tool updates is in Sec. 3.4
@@ 3.2 Inference
Plan generation. $\mathcal{D}$ is composed by $\mathcal{D}=\{\mathcal{D}_{p,s},\mathcal{D}_{p,f},\mathcal{D}_{c,s},\mathcal{D}_{c,f}\}$, where $\mathcal{D}_{p,s}$ and $\mathcal{D}_{p,f}$ contain correct and incorrect examples for plan generation respectively, and $\mathcal{D}_{c,s}$ and $\mathcal{D}_{c,f}$ contain correct and incorrect examples for program generations respectively. Given a task, we use the BERT model [20] to extract features of the given instruction and examples stored in $\mathcal{D}_{p,s}$ and $\mathcal{D}_{p,f}$. Then, we combine similar examples in $\mathcal{D}_{p,s}$ and $\mathcal{D}_{p,f}$ with the instruction to create a prompt. Finally, we send the prompt to LLMs to generate the plan in a one-go manner.
@@
Program generation. Similar to plan generation, we use LLMs to generate programs in a one-go manner. We select correct and incorrect examples of programs from $\mathcal{D}_{c,s}$ and $\mathcal{D}_{c,f}$. These examples are then combined with the plan as a prompt that is sent to LLMs for program generation.
@@
Tool execution. We utilize the interpreter module in [11] to parse the program, extracting tool names, inputs, and outputs of each step. CLOVR activates tools from a toolkit $\mathcal{T}$ that contains $15$ tools totally, including visual models and Python functions, as shown in Tab. 2.
@@ 3.3 Reflection
Result conversion. Since LLMs often struggle to identify errors by themselves [15, 56, 61], we provide the environmental feedback, desirable results, our wrong results, and intermediate results of each step for LLMs to better identify the error source. This requires us to convert visual results into textual form. For this purpose, we use the BLIP [26] model to convert images into languages by captioning.
Global reflection. CLOVA first uses global reflection to generate critiques in a one-go manner. The prompts are composed of task inputs, feedback on the task (e.g., desirable results in VQA tasks, or human comments on image editing tasks), generated plan and program, and intermediate results at each step. We send the prompts to LLMs to generate critiques that are used to update tools in the learning phase.
Local reflection. If CLOVA still fails after the tools are updated via global reflection and the learning phase –meaning the actual tools that lead to the faulty response are still to be found, we resort to local reflection to analyze each step of the program. Prompts are composed of the task inputs, feedback on the task, the steps that have been checked, and the current step that needs to be checked. Each step includes plans, programs, and intermediate results. We send the prompts to LLMs to infer whether this step has errors and reasons. Local reflection continues until an error location and reasons are identified for a step.
@@ 3.4 Learning
After obtaining tools that need to be updated from the reflection phase, CLOVA then moves to the learning phase to collect training data and goes through training-validation prompt tuning to update them, as shown in Fig. 5. Data collection. Since the tools that need to be updated can be rather different (a full list can be found in Tab. 2), we explore three manners to collect data in real-time: 1) We use LLMs to generate training data for the VQA tool. If reflection concludes that the VQA model makes errors, we combine the desirable response of the whole task and intermediate results of each step to prompt LLMs into inferring the correct output of the VQA tool. The question and the inferred output are then used to update the VQA tool. 2) We gather training data from open-vocabulary visual object grounding datasets (e.g., LVIS [10]) for the LOC and SEG tools. For example, if the reflection phase indicates that LOC does not work well for “horse”, CLOVA will select images and bounding boxes of horses from LVIS to update LOC. 3) We collect data by searching on the Internet for SELECT, CLASSIFY, and REPLACE tools. For instance, in case CLASSIFY is flagged by the reflection on not recognizing “horse” well, CLOVA will search for images of horses on the Internet to update CLASSIFY.

### v2

# Wenjuan Han3, Song-Chun Zhu1,2,4, Qing Li2 ✉

@@ Abstract
Utilizing large language models (LLMs) to compose off-the-shelf visual tools represents a promising avenue of research for developing robust visual assistants capable of addressing diverse visual tasks. However, these methods often overlook the potential for continual learning, typically by freezing the utilized tools, thus limiting their adaptation to environments requiring new knowledge. To tackle this challenge, we propose CLOVA, a Closed-LOop Visual Assistant, which operates within a framework encompassing inference, reflection, and learning phases. During the inference phase, LLMs generate programs and execute corresponding tools to complete assigned tasks. In the reflection phase, a multimodal global-local reflection scheme analyzes human feedback to determine which tools require updating. Lastly, the learning phase employs three flexible approaches to automatically gather training data and introduces a novel prompt tuning scheme to update the tools, allowing CLOVA to efficiently acquire new knowledge. Experimental findings demonstrate that CLOVA surpasses existing tool-usage methods by 5% in visual question answering and multiple-image reasoning, by 10% in knowledge tagging, and by 20% in image editing. These results underscore the significance of the continual learning capability in general visual assistants.
Figure 1: CLOVA is a general visual assistant that updates both LLMs and visual tools via inference, reflection, and learning in a closed-loop framework. During inference, CLOVA uses LLMs to integrate visual tools to accomplish given tasks. In reflection, CLOVA identifies tools that require updating based on human feedback. Finally, in learning, CLOVA collects data and updates the tools accordingly.
footnotetext: ✉  Corresponding author: Qing Li (dylan.liqing@gmail.com).
@@ 1 Introduction
The advancement of large language models (LLMs) [68, 47] alongside various visual tools (e.g., neural networks and OpenCV functions) [56, 57, 6, 45, 28] offers feasible avenues for constructing general visual assistants. When confronted with a task accompanied by language instructions, a common approach involves harnessing LLMs to generate programs, which are then executed using readily available visual tools to solve the task as dictated by the generated program [33, 10, 7, 41, 66, 61]. For instance, when posed with the query ”What is the person to the left of the umbrella doing?”, a viable solution entails LLMs sequentially performing the following steps: (1) utilizing a detection tool to locate the umbrella, (2) cropping the image region to the left of the umbrella, (3) using the detection tool to locate the person, and (4) finally querying a Visual Question Answering (VQA) tool with the question ”What is the person doing?”. Being highly compositional, such tool-usage methods demonstrate impressive performance and appealing explainability in tackling complex reasoning tasks, including VQA [31, 30, 11], mathematical reasoning [13, 14], and image editing [78, 79, 39, 73].
However, the potential for continual learning has been largely overlooked in existing tool-usage methods. Most of them simply freeze the used tools, which limits their applicability to environments where new knowledge is required, as depicted in Fig. 1. For instance, a user might instruct a visual assistant to label the face of the movie director Bong Joon-ho in a photograph. However, if the face recognition tool employed by the assistant fails to recognize Bong Joon-ho, it may provide an incorrect response. In such scenarios, it is expected that the assistant can learn this missing information about Bong Joon-ho and generalize it to other photographs. Thus, it is imperative to endow visual assistants with the learning capability, enabling them to swiftly acquire new knowledge from failures.
@@
In this paper, we propose CLOVA, a Closed-LOop Visual Assistant that updates used tools via closed-loop learning [32, 34] to better adapt to new environments, as illustrated in Figure 1. CLOVA consists of three phases: inference, reflection, and learning. During inference, CLOVA employs Large Language Models (LLMs) to generate programs and execute corresponding tools to accomplish the task at hand. Subsequently, in the reflection phase, CLOVA utilizes human feedback to provide critiques, identifying tools that require updates. Finally, in the learning phase, CLOVA autonomously collects data and updates tools accordingly. Thus, CLOVA facilitates autonomous tool updating, thereby continually enhancing their ability to adapt to diverse environments.
@@
To establish such a closed-loop learning framework, we must address three key challenges. Firstly, identifying tools that require updates is difficult due to the multi-step nature of generated programs and the diversity of errors within them. Secondly, automatically collecting training data is necessary as the knowledge to be learned is unpredictable. Thirdly, efficiently updating tools presents another obstacle, considering their scale and the quality of the collected data. Visual tools typically involve large neural networks, making them inefficient to update, and naive fine-tuning could result in unacceptable catastrophic forgetting [43]. Moreover, the presence of noise within the collected data further complicates the training process.
@@
We propose several techniques to tackle these challenges. First, we introduce a multimodal global-local reflection scheme, which resorts to LLMs to identify tools that need to be updated from both global and local aspects. For the second challenge, three data collection manners are employed, including inferring answers by LLMs, searching on the Internet, and searching from open-vocabulary datasets. Lastly, we develop a training-validation prompt tuning scheme for the tools, which includes instance-wise prompt tuning and a subsequent prompt validation stage, where learned prompts that fail to predict the validation data will be discarded. The learning phase also updates LLMs by storing correct examples and incorrect examples with critiques as in-context examples, which will be used in future inference. As a result, CLOVA efficiently updates tools in challenging environments with noisy data, while avoiding catastrophic forgetting.
@@
We apply CLOVA to compositional VQA and multiple-image reasoning tasks, using the GQA [19] and NLVRv2 [65] datasets. Additionally, we manually collect data for image editing and factual knowledge tagging tasks. CLOVA outperforms existing tool-usage methods by 5% in compositional VQA and multiple-image reasoning tasks, by 10% in knowledge tagging tasks, and by 20% in image editing tasks, showing the significance of the learning capability for general visual assistants.
In summary, our contributions are three-fold:
  • We build CLOVA, a visual assistant that updates its tools within a closed-loop learning framework for better adaptation to new environments.
  • We propose a multimodal global-local reflection scheme, capable of identifying tools in need of updates.
  • We employ three flexible manners to automatically collect training data and introduce a novel training-validation prompt tuning scheme to update tools efficiently while avoiding catastrophic forgetting.
@@ 2.1 General Visual Assistant
Benefiting from the advancements of LLMs [68, 47] and visual tools [56, 25, 45, 57], visual assistants have achieved great progresses. Some methods concatenate and train LLMs with visual tools in an end-to-end manner, where representative work includes LLaVA [37], Otter [26], MMICL [81], Kosmos-2 [52], and Flamingo [1], etc. In addition, some work extends the idea of tool usage for AI assistants from natural language processing [55, 49, 4, 54, 59, 12] to computer vision. By providing in-context examples, VISPROG [10] and ViperGPT [66] generate programs to use visual tools. Following this idea, some work improves performance by collecting instruction-following data [78, 38, 51], adding more tools [61, 35], and designing more dedicated tool-usage procedures [74, 79, 40, 41, 73, 39, 15]. The most related work to CLOVA is AssistGPT [7] and OpenAGI [8]. The two methods update LLMs after development through in-context learning and reinforcement learning, respectively. Different from them, CLOVA can update both LLMs and visual tools via its reflection and learning phases. This allows CLOVA to better adapt to new environments. In addition, the closed-loop framework enables us to set a separate training stage for tool-usage methods, going beyond zero-shot or few-shot visual assistants. Comparisons between CLOVA and some representative tool-usage methods are shown in Tab. 1.
Method Visual Tool Reflection Update LLMs Update VTs
@@
Table 1: Comparisons with representative tool-usage methods, where VTs means visual tools.
@@ 2.2 Reflection of LLMs
Reflection has become a remedy in case LLMs cannot generate good responses in a single attempt [48, 77, 50, 5]. Reflection methods send outputs back to LLMs to obtain critiques and further improve the outputs. These critiques take the form of scores or natural language [44, 63, 46, 84]. To generate better critiques, some methods employ instruction tuning [76, 58] or reinforcement learning [54, 3]. Recently, Huang et al. [17] revealed that LLMs struggle to provide accurate critiques for complex tasks. One way to address this issue is incorporating external information such as human-desired results into LLMs [77, 75, 2]. Unlike existing methods that rely solely on feedback in the language modality, our method generates reflection using all multimodal intermediate results. In addition, our method incorporates both global and local aspects for reflection, instead of only the global aspect. These bring more effective critiques for compositional tasks.
2.3 Prompt-based Learning
Prompt-based learning is an efficient technique to update neural networks, achieving impressive performance in both NLP [62, 21] and computer vision [85, 60, 86, 18]. Prompt engineering and prompt tuning are two kinds of commonly used methods. Prompt engineering develops interpretable tokens (e.g., texts and image regions) to guide model prediction, which are usually obtained by manually designing [67, 56], retrieval [80], and model generation [16, 53]. Prompt tuning learns vectors as prompts via gradient-based optimization. VPT [20] and CoOp [83] learn prompts for vision encoders and text encoders, respectively. To handle diverse data, CoCoOp [82] learns to generate prompts for unknown classes, ProDA [42] builds a Gaussian distribution for prompts, and MaPLe [23] learns both text and visual prompts. In addition, some methods employ prompts for continual learning, which learn prompts for different classes and produce adaptive prompts during inference. Representative methods include PIVOT [70], DualPrompt [71], and L2P [72]. Different from existing methods, our training-validation prompt tuning scheme discards harmful prompts, leading to more stable learning processes when the quality of training data is subpar.
@@ 3.1 Overview
As shown in Fig. 2, CLOVA has three phases: inference, reflection, and learning. In the inference phase, CLOVA uses LLMs to generate programs and executes corresponding tools to solve the task. The reflection phase introduces a multimodal global-local reflection scheme that uses LLMs to generate critiques, identifying which tool needs to be updated. During learning, we employ three manners to collect training data and use a training-validation prompt tuning scheme to update the tools.
@@
Table 2: Used tools in CLOVA, categorized based on whether the tool is updated in our method. Details of tool updates are in Sec. 3.4
@@ 3.2 Inference
Plan generation. The demonstration pool $\mathcal{D}$ is composed by $\mathcal{D}=\{\mathcal{D}_{p,s},\mathcal{D}_{p,f},\mathcal{D}_{c,s},\mathcal{D}_{c,f}\}$, where $\mathcal{D}_{p,s}$ and $\mathcal{D}_{p,f}$ contain correct and incorrect examples for plan generation respectively, and $\mathcal{D}_{c,s}$ and $\mathcal{D}_{c,f}$ contain correct and incorrect examples for program generation respectively. Given a task, we use the BERT model [22] to extract features of the given instruction and examples stored in $\mathcal{D}_{p,s}$ and $\mathcal{D}_{p,f}$. Then, we combine similar examples in $\mathcal{D}_{p,s}$ and $\mathcal{D}_{p,f}$ with the instruction to create a prompt. Finally, we send the prompt to LLMs to generate the plan in a one-go manner.
@@
Program generation. Similar to plan generation, we use LLMs to generate programs in a one-go manner. We select correct and incorrect examples of programs from $\mathcal{D}_{c,s}$ and $\mathcal{D}_{c,f}$. We combine these examples with the plan as a prompt and then send the prompt to LLMs for program generation.
@@
Tool execution. We utilize the interpreter module in [10] to parse the program, extracting tool names, inputs, and outputs of each step. CLOVR activates tools from a toolkit $\mathcal{T}$ that contains $15$ tools, including neural networks and Python functions, as shown in Tab. 2.
@@ 3.3 Reflection
Result conversion. Since LLMs often struggle to identify errors by themselves [17, 64, 69], we provide the human feedback, our wrong results, and intermediate results of each step for LLMs to better identify the error source. This requires us to convert visual results into textual form. For this purpose, we use the BLIP [28] model to convert images into languages by captioning.
Global reflection. CLOVA first uses global reflection to generate critiques in a one-go manner. The prompts are composed of task inputs, feedback on the task (e.g., desirable results in VQA tasks, or human comments in image editing tasks), generated plans and programs, and intermediate results at each step. We send the prompts to LLMs to generate critiques that are used to update tools in the learning phase.
Local reflection. If CLOVA still fails after the tools are updated via global reflection and the learning phase–meaning the actual tools that lead to the faulty response are still to be found, we resort to local reflection to analyze each step of the program. Prompts are composed of the task inputs, feedback on the task, the steps that have been checked, and the current step that needs to be checked. Each step includes plans, programs, and intermediate results. We send the prompts to LLMs to infer whether this step has errors and the reasons. Local reflection continues until an error location and reasons are identified for a step.
@@ 3.4 Learning
3.4.1 Updating tools with prompt tuning
After identifying tools that need to be updated from the reflection phase, CLOVA then moves to the learning phase to collect training data and goes through training-validation prompt tuning to update the tools, as shown in Fig. 5.
Data collection. Since the tools that need to be updated can be rather different (a full list can be found in Tab. 2), we explore three manners to collect data online. (1) We use LLMs to generate training data for the VQA tool. If reflection concludes that the VQA tool makes errors, we combine the desirable response of the whole task and intermediate results of each step to prompt LLMs into inferring the correct output of the VQA tool. The question and the inferred output are then used to update the VQA tool. (2) We gather training data from open-vocabulary visual object grounding datasets (e.g., LVIS [9]) for the LOC and SEG tools. For example, if the reflection phase indicates that LOC does not work well for the visual concept “horse”, CLOVA will select images and bounding boxes of horses from LVIS to update LOC. (3) We collect data by searching on the Internet for the SELECT, CLASSIFY, and REPLACE tools. For instance, If CLASSIFY is marked as unable to recognize “horse” during the reflection phase, CLOVA will search the Internet for images of horses to update CLASSIFY.